# Blending Data-Driven Priors in Dynamic Games

### 21 Feb 2024 v1

@@ II-A Game-Theoretic Motion Planning
Multi-modality in planning can arise from uncertain or unknown objectives. For example, in a merge setting, which player prefers to go first may directly alter the outcome of interaction, and this objective uncertainty may lead to indecision. Cleac’h et al. [18] introduce a planning for uncertain linear cost models by estimating cost parameters using an unscented Kalman filter. Hu et al. [35] study emergent coordination in multi-agent interactions when cost parameters follow an opinion-dynamic interaction over a graph. Recent work [36] uses differentiable optimization to learn open-loop mixed strategies for non-cooperative games based on trajectory data. Liu et al. [17] extend this idea to compute generalized equilibria, a setting where hard constraints are present, and use deep learning to bolster the computation performance of the game solver. Another follow-up work [37] further studies the setting of learning a feedback Nash equilibrium strategy.
To the best of our knowledge, our approach is the first model-based game-theoretic planner that can incorporate multi-modal priors from data-driven trajectory prediction models for guided exploration and producing scene-consistent interactions.
@@ IV-C1 The LQL approximation
Backward Pass. During the backward pass, we linearize dynamics (1) and quadraticize cost (2) around ${\eta}$ to obtain a linear dynamical system and quadratic cost functions that fit into the assumption of KL-LQG game (c.f. Theorem 1). Next, we use the Laplace approximation [84, Chapter 4] to obtain a Gaussian distribution that captures the reference policy locally around ${\eta}$. Specifically, the conditional probability distribution of a reference policy ${\tilde{\pi}}^{i}_{t}({u}^{i}_{t}|{x}_{t})$ is approximated as:
@@
Line Search. A common issue in ILQ-based trajectory optimization is that directly applying the mean of the KL-LQG strategy in the forward pass may cause the system to deviate too much from the nominal trajectory ${\eta}$, resulting in divergence of the algorithm. A possible remedy is line search [85, Chapter 3], which controls the trajectory update by scaling down the KL-LQG strategy parameters. Specifically, for a line search parameter $\varepsilon\in(0,1]$ and KL-LQG strategy $\mathcal{N}(-K^{i}_{t}{x}_{t}-\kappa^{i}_{t},\Sigma_{t}^{i})$, the adjusted strategy is $\mathcal{N}(-\varepsilon K^{i}_{t}{x}_{t}-\varepsilon\kappa^{i}_{t},\varepsilon\Sigma_{t}^{i})$, which is then used in the forward pass to obtain a new nominal trajectory ${\eta}$. The procedure starts with $\varepsilon\leftarrow 1$ and repeats by decreasing $\varepsilon$ by half until a convergence criterion is met. Two commonly used criteria are decreasing of the social cost (i.e., the sum of all players’ costs) and that the updated nominal trajectory is sufficiently close to the previous one.
@@ V-B Autonomous Car Racing
Multi-modal KLGame. Motivated by the result shown above for Basic KLGame, where blending only an aggressive reference policy can lead to unsafe outcomes, this experiment demonstrates how a multi-modal KLGame (MM-KLGame) (c.f. Sec. IV-C2) with an additional safety-enhancing mode performs in racing scenario. The second mode we use is a following policy that controls the ego car to (i) follow the rival while keeping a constant distance when the rival is leading or (ii) defend against the rival after overtaking it. During each planning cycle, instead of picking the mode via sampling, we use rollout-based safety filters [91, Section 3.3] and perform a collision check for the KLGame-optimized trajectories associated with the overtaking mode: if they are collision-free, we choose the overtaking mode; otherwise, we pick the following mode. Again, we vary the value of $\lambda$ and inspect the qualitative difference in the resulting ego behaviors. The simulation snapshots are shown in Fig. 8. When $\lambda$ takes a relatively small value, $\lambda=5$, the MM-KLGame is not sufficiently guided by the overtaking policy and fails to overtake the rival. When $\lambda=50$, the ego using MM-KLGame behaves similarly to the unimodal case and safely overtakes the rival. However, in the case when $\lambda=5\times 10^{4}$, instead of incurring a collision as in the unimodal KLGame case, the MM-KLGame leads to a safe overtake thanks to the additional following mode in the multi-modal reference policy. We also apply the reference policy to the ego using the same mode selection rule. Interestingly, the ego car fails to overtake the rival due to a more frequent application of the following policy triggered by the unsafe overtaking policy, which ultimately slows down the ego vehicle. This example also demonstrates the potential of KLGame using a multi-modal reference policy that mixes a data-driven policy and an optimization-based policy.

### 23 Feb 2024 v2

@@ II-A Game-Theoretic Motion Planning
Multi-modality in planning can arise from uncertain or unknown objectives. For example, in a merge setting, which player prefers to go first may directly alter the outcome of interaction, and this objective uncertainty may lead to indecision. Cleac’h et al. [18] introduce a planning for uncertain linear cost models by estimating cost parameters using an unscented Kalman filter. Chen et al. [35] introduce an MPC-based framework for interactive trajectory optimization using deep-learning predictions as input. Hu et al. [36] study emergent coordination in multi-agent interactions when cost parameters follow an opinion-dynamic interaction over a graph. Recent work [37] uses differentiable optimization to learn open-loop mixed strategies for non-cooperative games based on trajectory data. Liu et al. [17] extend this idea to compute generalized equilibria, a setting where hard constraints are present, and use deep learning to bolster the computation performance of the game solver. Another follow-up work [38] further studies the setting of learning a feedback Nash equilibrium strategy. In the potential game setting, Diehl et al. [39] show that multiple open-loop Nash equilibria can be computed in parallel using deep-learned game parameters.
To the best of our knowledge, our approach is the first closed-loop, model-based game-theoretic planner that can incorporate multi-modal priors from data-driven trajectory prediction models for guided exploration and producing scene-consistent interactions. Our method makes only mild assumptions on the environment dynamics and planning costs (e.g. being differentiable). Moreover, our provision of a scalar regularization parameter permits tunable adherence to data-driven behavior at inference time, in contrast to traditional game-theoretic planning and end-to-end equilibrium learning, which are rigid in their respective assumptions about planning objectives (i.e. task-driven or data-driven).
@@ IV-C1 The LQL approximation
Backward Pass. During the backward pass, we linearize dynamics (1) and quadraticize cost (2) around ${\eta}$ to obtain a linear dynamical system and quadratic cost functions that fit into the assumption of KL-LQG game (c.f. Theorem 1). Next, we use the Laplace approximation [85, Chapter 4] to obtain a Gaussian distribution that captures the reference policy locally around ${\eta}$. Specifically, the conditional probability distribution of a reference policy ${\tilde{\pi}}^{i}_{t}({u}^{i}_{t}|{x}_{t})$ is approximated as:
@@
Line Search. A common issue in ILQ-based trajectory optimization is that directly applying the mean of the KL-LQG strategy in the forward pass may cause the system to deviate too much from the nominal trajectory ${\eta}$, resulting in divergence of the algorithm. A possible remedy is line search [86, Chapter 3], which controls the trajectory update by scaling down the KL-LQG strategy parameters. Specifically, for a line search parameter $\varepsilon\in(0,1]$ and KL-LQG strategy $\mathcal{N}(-K^{i}_{t}{x}_{t}-\kappa^{i}_{t},\Sigma_{t}^{i})$, the adjusted strategy is $\mathcal{N}(-\varepsilon K^{i}_{t}{x}_{t}-\varepsilon\kappa^{i}_{t},\varepsilon\Sigma_{t}^{i})$, which is then used in the forward pass to obtain a new nominal trajectory ${\eta}$. The procedure starts with $\varepsilon\leftarrow 1$ and repeats by decreasing $\varepsilon$ by half until a convergence criterion is met. Two commonly used criteria are decreasing of the social cost (i.e., the sum of all players’ costs) and that the updated nominal trajectory is sufficiently close to the previous one.
@@ V-B Autonomous Car Racing
Multi-modal KLGame. Motivated by the result shown above for Basic KLGame, where blending only an aggressive reference policy can lead to unsafe outcomes, this experiment demonstrates how a multi-modal KLGame (MM-KLGame) (c.f. Sec. IV-C2) with an additional safety-enhancing mode performs in racing scenario. The second mode we use is a following policy that controls the ego car to (i) follow the rival while keeping a constant distance when the rival is leading or (ii) defend against the rival after overtaking it. During each planning cycle, instead of picking the mode via sampling, we use rollout-based safety filters [92, Section 3.3] and perform a collision check for the KLGame-optimized trajectories associated with the overtaking mode: if they are collision-free, we choose the overtaking mode; otherwise, we pick the following mode. Again, we vary the value of $\lambda$ and inspect the qualitative difference in the resulting ego behaviors. The simulation snapshots are shown in Fig. 8. When $\lambda$ takes a relatively small value, $\lambda=5$, the MM-KLGame is not sufficiently guided by the overtaking policy and fails to overtake the rival. When $\lambda=50$, the ego using MM-KLGame behaves similarly to the unimodal case and safely overtakes the rival. However, in the case when $\lambda=5\times 10^{4}$, instead of incurring a collision as in the unimodal KLGame case, the MM-KLGame leads to a safe overtake thanks to the additional following mode in the multi-modal reference policy. We also apply the reference policy to the ego using the same mode selection rule. Interestingly, the ego car fails to overtake the rival due to a more frequent application of the following policy triggered by the unsafe overtaking policy, which ultimately slows down the ego vehicle. This example also demonstrates the potential of KLGame using a multi-modal reference policy that mixes a data-driven policy and an optimization-based policy.